# Calibration Attack: A Framework For Adversarial Attacks Targeting Calibration

### v1

@@ Appendix D Calibration Metric Formulation
Here we forumalate the two calibration metrics that we use in our experiments. As Equation 2 is an idealized representation of miscalibration that is intractable to calculate, approximations have been developed which are grouped into the more common binning-based metrics, and non-binning based metrics.
@@
Appendix E Adversarial Attack Detection Details
Local Intrinsic Dimensionality (LID) (Ma et al., 2018). This detection method exploits the estimated Local Intrinsic Dimensionality (LID) characteristics across different layers of a model of a set of adversarial examples, which are found to be notably different than that of clean samples or those with added random noise. First, a training set is made up of clean, noisy, and adversarial examples, and a simple classifier (logistic regression) is trained to discriminate between adversarial and non-adversarial samples. For each training minibatch, the input features to the classifier are generated based on the estimated LID across different layers for all of the samples. The two main hyperparameters for this method are batch size and the number of nearest neighbours involved in estimating the LID. We choose a consistent batch size of 100 in line with previous work such as (Harder et al., 2021), and for each case we test the possible nearest neighbors from the following list $\{10,20,30,40,50,60,70,80,90\}$ and report the results for the best value, which vary for different datasets and models. We use the implementation of LID from (Lee et al., 2018).
Mahalanobis Distance (MD) (Lee et al., 2018). The premise behind this method is to use a set of training samples to fit a class-conditional Gaussian distribution based on the empirical class means and empirical covariance of the training samples. Given a test sample, the Mahalanobis distance with respect to the closest class-conditional distribution is found and taken as the confidence score. A logistic regression detector is built from this which determines whether a sample is adversarial. The main hyperparameter for this method is the magnitude of the noise used, which we vary between $\{0.0,0.01,0.005,0.002,0.0014,0.001,0.0005\}$ for each case and pick the value that results in the highest detection accuracy. In addition, calculating the mean and covariance is necessary to use the method, which we utilize the respective training set to do for each dataset. We use the implementation of MD from (Lee et al., 2018).
@@ Appendix E Adversarial Attack Detection Details
Appendix F Additional Analysis and Results
@@ F.1 White Box Variation of Attack
F.2 Detailed Analysis of Epsilon and Iterations
Epsilon. The $\epsilon$ parameter plays a major role in adversarial attacks, as it controls how much noise can be added when creating perturbations. Although setting a higher $\epsilon$ value for an attack lets it easier and more efficient for the algorithm to create adversarial examples, it potentially cause the visual changes to images more perceptible, so a small $\epsilon$ is preferable while still being able to produce good adversarial examples. In the case of calibration attack, there is no need to go as far as flipping a label, so lower $\epsilon$-bounds have the potential to create some miscalibration. To provide further details on our results in Figure 2, for the leftmost figure as mentioned previous we tested on CIFAR-100 using ResNet-50. The five different $\epsilon$ values we use are (0.005, 0.01, 0.05, 0.1, 0.25) after being attacked using all four of the attacks with the other settings the same as in Appendix  B, with the results averaged over three models. In addition to the miscalibration being strong for most of the attacks at low $\epsilon$ values, we can see that maximum miscalibration attack consistently outperforms the rest across the different values. The underconfidence attack does not have much change with higher $\epsilon$, but it is largely because the models has already almost reached the peak level of attacking the confidence with low epsilon values, and as such does not have a large effect on ECE. As middle figure largely displays the same trends as ResNet, revealing that the results are not architecture dependant. The rightmost figure uses ResNet and goes over the same $\epsilon$ values as before, except the maximum miscalibration attack is run over both Caltech-101 and GTSRB models. Again the trends are similar, although the increase in ECE is not as severe as for CIFAR-100.
@@ F.2 Detailed Analysis of Epsilon and Iterations
Iterations. Expanding on the results in the rightmost figure in Figure 2, the number of iterations of the maximum miscalibration attack is varied from 100, 500, 1000, to 5000, whilst attack both ViT and ResNet models trained and tested on CIFAR-100 with the same settings as in Appendix  B. We note how the ECE begins to saturate at close to 500 iterations, after which the benefits of running the attack longer are minor, though even at 100 iterations the ResNet model becomes heavily miscalibrated despite the standard $\epsilon$ value of 0.05 being used. In our tests showing the effectiveness of the original SA versus its calibration attack version, seen in Figure 4 we specifically compare over accuracy and ECE between the maximum miscalibration attack and the regular untargeted Square Attack across the four aforementioned iteration values for both ResNet and ViT on CIFAR-100. As expected, Square Attack greatly reduces the accuracy even with a small number of iterations. Nevertheless, in terms of ECE, the calibration attacks consistently produce higher amounts of miscalibration compared to the original Square Attack across the different iteration amounts.
F.3 t-SNE Visualizations
@@
To help visualize the effect of each of the attack types in latent space and to confirm they are having the expected effects, we run a t-SNE analysis (van der Maaten & Hinton, 2008) on the representations of ResNet-50 right before the classification layer. The datasets we use throughout this study, with their large number of classes, are not ideal for visualization purposes. Instead, we create a binary subset using CIFAR-100 by taking all of the images from two arbitrary classes, bicycles and trains. We create a separate training set and test set to perform this procedure independently, and fine-tune a ResNet-50 model on the training set. The specific details are similar to those described in Section C for CIFAR-100 ResNet. We train the model for 5 epochs with a learning rate of 0.005. The attack settings are the same as in Section B for the $l_{\infty}$ version, and we only run the attacks for 500 iterations. We run the t-SNE analysis on a balanced slice of 200 images from the new test set for easy visualization purposes, before and after all of the different attacks. The model achieves 95% accuracy on the full test set. Figure 5 shows the graphs. It can be seen the effect on the representations for the adversarially attacked data is as expected. The overconfidence attack causes the representations for both class predictions, even incorrect ones, to be split apart as much as possible, while the underconfidence attack causes a more jumbled representation between the two classes with most falling closely to the decision boundary. The maximum miscalibration attack has a similar effect to the underconfidence attack, except the misclassified images are pushed far away from the decision boundary to make it appear as if the model is more confident in its decisions. Lastly, the random attack causes two distinct random clusters for each prediction type to form, as random data points are pushed to be more overconfident or more underconfident than they originally were. With these results, we can see visually confirm that the attacks possess their intended behaviour.
F.4 Imbalance Ratio
@@
Dataset imbalance has a profound effect on how a model learns and how well it performs, with detrimental effects occurring when imbalance ratios are very high. With how common imbalanced data distributions are in real world scenarios, we believe it is worth studying the influence of imbalance ratio and its relationship with robustness against calibration attacks as an additional point of analysis. We choose CIFAR-100 as our primary dataset for this analysis, and we follow the procedures in Tang et al. (2020) and Cao et al. (2019) to create training sets with long-tail imbalance. This is a form of imbalance where the sample sizes in the classes follow an exponential decay. We use the variable $\rho$ to denote the ratio between sample sizes of the class with the smallest sample size, and those of the one chosen to be the biggest. We create training sets with $\rho$ values of 0.01, 0.02, and 0.01 (for 1:100, 1:50 and 1:10 ratios of smallest to biggest class). We then train 3 ResNet-50 and 3 ViT models on each imbalanced set. The training details are again the same as those described in the general settings Section C, although the training times are different. 15 epochs is used to training the 1:100 ratio models, while 10 epochs is used for the rest. We subject the models to the maximum miscalibration attack using the same settings as in Section B for CIFAR-100 (test data is balanced), and calculate the resulting average and deviation of the pre and post attack ECE, average number of queries, and accuracy. The graphs displaying the results can be seen in Figure 6. Unsurprisingly, the higher the imbalance ratio, the lower the accuracy is on the balanced data. In terms of robustness, the more balanced the data the more resistant it is against getting miscalibrated from the attacks, for both the ResNet and ViT architectures. This is similar to the trends in the inherent miscalibration present before the attacks, although the calibration differences between the different ratio models are not as severe, and ViT at the 1:500 imbalance ratio is the best calibrated beforehand but becomes the worst after the attack. The trend in the number of queries it takes for a successful attack is reversed for ResNet and ViT, with ViT requiring more queries the more balanced the data is, while ResNet is vice-versa. Overall, dataset imbalance does not create favourable conditions for robustness, though the use of imbalance data techniques could potentially remedy some of these issues.
F.5 GradCAM Visualization Details
@@
Given the effectiveness of the attacks at leading a model to produce highly miscalibrated outputs, for both base styles of attacks, we endeavour to explore whether they also lead to any changes in where the model focuses on in an image when making its decision, and especially with novel overconfidence attack. Knowing this can lead to further insights as to how models are affected by various forms of the attacks. To accomplish this analysis, we use Grad-CAM (Selvaraju et al., 2017), a popular visualization method that produces a coarse localization map highlighting the most important regions in an image that the model uses when making its prediction by making use of the gradients from the final convolutional layer (or a specific layer of choice) of a network. We apply Grad-CAM to our ResNet-50 models fine-tuned on Caltech-101 to images from the Caltech-101 test set before and after the underconfidence and overconfidence attacks at the standard attack settings used in Section B and using the GradCAM implementation of Gildenblat & contributors (2021). Since the method calculates relative to a specific class, we do so in-terms of the predicted class. Figure 3 shows the results with some representative images. We specifically choose images where the attacks led to large change in predicted confidence (at least 10%). On the whole, we have observed that the coarse localization maps have minimal to no noticeable changes after the adversarial images are produced, especially in the case of the overconfidence attacked images. This leads us to believe the primary mechanism of the attacks changing the model confidence is in the final classification layer as opposed to the convolutional layers. This analysis also shows that it might be difficult to identify these attacks are occurring based on these types of gradient visualization methods.
F.6 $L_{2}$ Results
Table 6: Results of the $l_{2}$ calibration attacks for the three different datasets.

### v2

@@ Appendix E Calibration Metric Formulation
Here we formulate the two calibration metrics that we use in our experiments. As Equation 2 is an idealized representation of miscalibration that is intractable, approximations have been developed which are grouped into the more common binning-based metrics, and non-binning based metrics.
@@
Appendix F Adversarial Attack Detection Details
Local Intrinsic Dimensionality (LID) (Ma et al., 2018). This detection method exploits the estimated Local Intrinsic Dimensionality (LID) characteristics across different layers of a model of a set of adversarial examples, which are found to be notably different than that of clean datapoints or those with added random noise. First, a training set is made up of clean, noisy, and adversarial examples, and a simple classifier (logistic regression) is trained to discriminate between adversarial and non-adversarial examples. For each training minibatch, the input features to the classifier are generated based on the estimated LID across different layers for all of the datapoints. The hyperparameters for this method are batch size and the number of nearest neighbours involved in estimating the LID. We choose a consistent batch size of 100 in line with previous work such as (Harder et al., 2021), and for each case we test the possible nearest neighbors from the following list $\{10,20,30,40,50,60,70,80,90\}$ and report the results for the best value, which vary for different datasets and models. We use the implementation from Lee et al. (2018).
Mahalanobis Distance (MD) (Lee et al., 2018). The premise behind this method is to use a set of training datapoints to fit a class-conditional Gaussian distribution based on the empirical class means and empirical covariance of the training datapoints. Given a test datapoint, the Mahalanobis distance with respect to the closest class-conditional distribution is found and taken as the confidence score. A logistic regression detector is built from this which determines whether a datapoint is adversarial. The main hyperparameter for this method is the magnitude of the noise used, which we vary between $\{0.0,0.01,0.005,0.002,0.0014,0.001,0.0005\}$ for each case and pick the value that results in the highest detection accuracy. In addition, calculating the mean and covariance is necessary to use the method, which we utilize the respective training set to do for each dataset. We use the implementation of MD from (Lee et al., 2018).
@@ Appendix F Adversarial Attack Detection Details
Appendix G Additional Analysis and Results
@@ G.1 White-box Calibration Attack
G.2 Detailed Setup and Comparison of Efficiency of Underconfidence vs. Overconfidence Attacks.
To understand which form of attack is most query efficient when the amount of change in confidence is the same, for each attack type we identify all of images in the test set that are around a given confidence level. We use the corresponding attack to made the model produce either an increase of 10% in confidence, or a decrease of 10%. We choose two base confidence levels of 80% and 90% and find all the data points within 1% of each. When an attack causes a change at or past the set threshold for the given goal probability, the attack stops and the number of queries is recorded. The results can be seen in Table 3. The consistent pattern we observe for both base confidence levels is that it takes notably fewer queries to create underconfidence than overconfidence, and the former attack is more effective at affecting the average confidence.
G.3 Detailed Analysis of Epsilon and Iterations
Epsilon. The $\epsilon$ parameter plays a major role in adversarial attacks, as it controls how much noise can be added when creating perturbations. Although setting a higher $\epsilon$ value for an attack lets it easier and more efficient for the algorithm to create adversarial examples, it potentially cause the visual changes to images more perceptible, so a small $\epsilon$ is preferable while still being able to produce good adversarial examples. In the case of calibration attack, there is no need to go as far as flipping a label, so lower $\epsilon$-bounds have the potential to create some miscalibration. To provide further details on our results in Figure 2, for the leftmost figure as mentioned previous we tested on CIFAR-100 using ResNet-50. The five different $\epsilon$ values we use are (0.005, 0.01, 0.05, 0.1, 0.25) after being attacked using all four of the attacks with the other settings the same as in Appendix  C, with the results averaged over three models. In addition to the miscalibration being strong for most of the attacks at low $\epsilon$ values, we can see that maximum miscalibration attack consistently outperforms the rest across the different values. The underconfidence attack does not have much change with higher $\epsilon$, but it is largely because the models has already almost reached the peak level of attacking the confidence with low epsilon values, and as such does not have a large effect on ECE. As middle figure largely displays the same trends as ResNet, revealing that the results are not architecture dependant. The rightmost figure uses ResNet and goes over the same $\epsilon$ values as before, except the maximum miscalibration attack is run over both Caltech-101 and GTSRB models. Again the trends are similar, although the increase in ECE is not as severe as for CIFAR-100.
@@ G.3 Detailed Analysis of Epsilon and Iterations
Iterations. Expanding on the results in the rightmost figure in Figure 2, the number of iterations of the maximum miscalibration attack is varied from 100, 500, 1000, to 5000, whilst attack both ViT and ResNet models trained and tested on CIFAR-100 with the same settings as in Appendix  C. We note how the ECE begins to saturate at close to 500 iterations, after which the benefits of running the attack longer are minor, though even at 100 iterations the ResNet model becomes heavily miscalibrated despite the standard $\epsilon$ value of 0.05 being used. In our tests showing the effectiveness of the original SA versus its calibration attack version, seen in Figure 4 we specifically compare over accuracy and ECE between the maximum miscalibration attack and the regular untargeted Square Attack across the four aforementioned iteration values for both ResNet and ViT on CIFAR-100. As expected, Square Attack greatly reduces the accuracy even with a small number of iterations. Nevertheless, in terms of ECE, the calibration attacks consistently produce higher amounts of miscalibration compared to the original Square Attack across the different iteration amounts.
G.4 t-SNE Visualizations
@@
To help visualize the effect of each of the attack types in latent space and to confirm they are having the expected effects, we run a t-SNE analysis (van der Maaten & Hinton, 2008) on the representations of ResNet-50 right before the classification layer. The datasets we use throughout this study, with their large number of classes, are not ideal for visualization purposes. Instead, we create a binary subset using CIFAR-100 by taking all of the images from two arbitrary classes, bicycles and trains. We create a separate training set and test set to perform this procedure independently, and fine-tune a ResNet-50 model on the training set. The specific details are similar to those described in Section D for CIFAR-100 ResNet. We train the model for 5 epochs with a learning rate of 0.005. The attack settings are the same as in Section C for the $l_{\infty}$ version, and we only run the attacks for 500 iterations. We run the t-SNE analysis on a balanced slice of 200 images from the new test set for easy visualization purposes, before and after all of the different attacks. The model achieves 95% accuracy on the full test set. Figure 5 shows the graphs. It can be seen the effect on the representations for the adversarially attacked data is as expected. The overconfidence attack causes the representations for both class predictions, even incorrect ones, to be split apart as much as possible, while the underconfidence attack causes a more jumbled representation between the two classes with most falling closely to the decision boundary. The maximum miscalibration attack has a similar effect to the underconfidence attack, except the misclassified images are pushed far away from the decision boundary to make it appear as if the model is more confident in its decisions. Lastly, the random attack causes two distinct random clusters for each prediction type to form, as random data points are pushed to be more overconfident or more underconfident than they originally were. With these results, we can see visually confirm that the attacks possess their intended behaviour.
G.5 Imbalance Ratio
@@
Dataset imbalance has a profound effect on how a model learns and how well it performs, with detrimental effects occurring when imbalance ratios are very high. With how common imbalanced data distributions are in real world scenarios, we believe it is worth studying the influence of imbalance ratio and its relationship with robustness against calibration attacks as an additional point of analysis. We choose CIFAR-100 as our primary dataset for this analysis, and we follow the procedures in Tang et al. (2020) and Cao et al. (2019) to create training sets with long-tail imbalance. This is a form of imbalance where the datapoint sizes in the classes follow an exponential decay. We use the variable $\rho$ to denote the ratio between datapoint sizes of the class with the smallest datapoint size, and those of the one chosen to be the biggest. We create training sets with $\rho$ values of 0.01, 0.02, and 0.01 (for 1:100, 1:50 and 1:10 ratios of smallest to biggest class). We then train 3 ResNet-50 and 3 ViT models on each imbalanced set. The training details are again the same as those described in the general settings Section D, although the training times are different. 15 epochs is used to training the 1:100 ratio models, while 10 epochs is used for the rest. We subject the models to the maximum miscalibration attack using the same settings as in Section C for CIFAR-100 (test data is balanced), and calculate the resulting average and deviation of the pre and post attack ECE, average number of queries, and accuracy. The graphs displaying the results can be seen in Figure 6. Unsurprisingly, the higher the imbalance ratio, the lower the accuracy is on the balanced data. In terms of robustness, the more balanced the data the more resistant it is against getting miscalibrated from the attacks, for both the ResNet and ViT architectures. This is similar to the trends in the inherent miscalibration present before the attacks, although the calibration differences between the different ratio models are not as severe, and ViT at the 1:500 imbalance ratio is the best calibrated beforehand but becomes the worst after the attack. The trend in the number of queries it takes for a successful attack is reversed for ResNet and ViT, with ViT requiring more queries the more balanced the data is, while ResNet is vice-versa. Overall, dataset imbalance does not create favourable conditions for robustness, though the use of imbalance data techniques could potentially remedy some of these issues.
G.6 GradCAM Visualization Details
@@
Given the effectiveness of the attacks at leading a model to produce highly miscalibrated outputs, for both base styles of attacks, we endeavour to explore whether they also lead to any changes in where the model focuses on in an image when making its decision, and especially with novel overconfidence attack. Knowing this can lead to further insights as to how models are affected by various forms of the attacks. To accomplish this analysis, we use GradCAM (Selvaraju et al., 2017), a popular visualization method that produces a coarse localization map highlighting the most important regions in an image that the model uses when making its prediction by making use of the gradients from the final convolutional layer (or a specific layer of choice) of a network. We apply GradCAM to our ResNet-50 models fine-tuned on Caltech-101 to images from the Caltech-101 test set before and after the underconfidence and overconfidence attacks at the standard attack settings used in Section C and using the GradCAM implementation of Gildenblat & contributors (2021). Since the method calculates relative to a specific class, we do so in-terms of the predicted class. Figure 3 shows the results with some representative images. We specifically choose images where the attacks led to large change in predicted confidence (at least 10%). On the whole, we have observed that the coarse localization maps have minimal to no noticeable changes after the adversarial images are produced, especially in the case of the overconfidence attacked images. This leads us to believe the primary mechanism of the attacks changing the model confidence is in the final classification layer as opposed to the convolutional layers. This analysis also shows that it might be difficult to identify these attacks are occurring based on these types of gradient visualization methods.
G.7 Certified Confidence Scores
In this section, we follow recent work in the domain of providing provable guarantees on the robustness of confidence scores by examining the effect on a smoothed classifier and its bounds when handling calibration attacked data, particularly in the case of overconfidence attacks. Given the lack of label flipping, we expect that adversarial examples generated by calibration attack lie close to the original in the data manifold, thereby having a minimum effect on the certified confidence. To test this hypothesis, we closely follow Kumar et al. (2020) on certification to provide lower bounds on the confidence of a smoothed classifier. A strong lower bound can be produced by using the probability distribution of the confidence scores of a Gaussian cloud around the input image using Neyman-Pearson lemma to calculate this for a given certified radius. We select a sub-sample of 100 random datapoints from CIFAR-100 and use our base ResNet-50 models that is attacked using the underconfidence and overconfidence attacks using our standard settings, and average the expected confidence score lower bound produced by the smoothed model across different Radii. We utilize the version of the method that uses CDF information and use a smoothing faction value of $\sigma=0.25$, failure probability $\alpha=0.001$ and use 100,000 Monte-Carlo datapoints for the estimation. The results for the certified smoothed model cab be seen in Figure 7. We observe that the average lower bound for the underconfidence attack is lower than the base version and overconfidence attack, although it is not dramatically different due to the huge error bounds. Nevertheless, it appears that the smoothed model works well at being robust, particularly on the overconfidence attacked datapoints, but despite the expected confidence, the average certified accuracy for each set of data each was 74%, 59%, and 76% respectively, so the smoothed classifier struggles at performing more accurately on the perturbed datapoints. Overall, this method of guarantee could be a useful tool for counteracting calibration attacks without using specific defences, particularly when using classifiers for safety-critical tasks that rely on confidence scores.
Figure 7: Expected Confidence Score lower bounds for a set of CIFAR-100 images (original and attacked using the underconfidence and overconfidence attacks) based on a ResNet-50 classifier.
G.8 Detailed Results and Analyses on Recalibration and Defence
Table 4 shows the experiment results of query efficiency, accuracy, along with the ECE and KS error, before (PrECE and PrKS) and after the attack (PrECE and PrKS).
The RobustBench models compromise substantially on accuracy, and have a high level of miscalibration on clean data. They do largely avoid getting extremely miscalibrated as a result of the attacks compared to the defenceless models, except the top model on the leaderboard. Nevertheless, their high inherent miscalibration means they are unfavourable in situations where the model must be well calibrated.
In terms of the calibration methods, TS tends to be among the best methods at reducing calibration error prior to the attacks, but after the attacks it offers very little benefit compared to the vanilla models. The MD-TS method is similar with its prior calibration being solid, but post attack it only brings minor benefits over TS in most cases. It appears that this is due in part to not finding the ideal temperature parameter for each image due to the difficulty of identifying the correct image domains, as in, recognizing which form of attack occurred. The Splines method is similar in its pre-attack calibration benefits to TS, but differs greatly in its performance post-attack. In some cases, like CIFAR-100 ResNet, it is easily the worst performing defence method. In other cases, particularly for Caltech-101 and GTSRB ViT, it is able to keep ECE at relatively reasonable values post-attack. This discrepancy shows that finding an ideal recalibration function has the potential to be a strong defence. The training-based DCA and SAM methods tend to bring few benefits after being attacked, even when they improve the calibration on clean data, the post-attack ECE and KS errors are not substantially different compared to the vanilla models.
The performance of the regular adversarial defence techniques is mixed. For robustness, AAA in most cases achieves the lowest post-attack ECE. Even in the best cases like Caltech-101 ResNet, ECE tends to be at least double compared pre-attack, and in most cases we still observed multiple-fold increases. This technique is also among the poorest calibrated on clean data. Regarding AT, our approach does not compromise on accuracy and miscalibration on clean data. It brings notable robustness, especially compared to the calibration methods, but it is not among the strongest.
Lastly, CS is the strongest methods at maintaining low calibration error on post-attack datapoints. Moreover, the technique tends to have better calibration error on clean data compared to AAA. It shows how it is key that high confidence values are retained to have decent calibration after the attacks. Altogether, despite some promising results with the defences, as a whole there are still limitations particularly with the strongest adversarial defences. The compromise of poor ECE on clean data for better calibration robustness against the attacks that we observe, as well as the general inconsistent performance means that further refinement on defences is warranted.
G.9 Results of $l_{2}$ calibration attacks
Table 6: Results of the $l_{2}$-based calibration attacks on three datasets.